OpenStudy (anonymous):

Help me with finding the limit of this sequence again, I have to use the squeeze theorem here.

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} n ^{n}/n!\]

OpenStudy (anonymous):

It would seem that the limit is infinity, right?

OpenStudy (anonymous):

So to use squeeze theorem, we just need something which is obviously less than that function, yet approaches infinity.

OpenStudy (anonymous):

Can u please elaborate?

OpenStudy (anonymous):

On why I think it approaches infinity?

OpenStudy (anonymous):

The basic intuition is this that it approaches to infinity but to use squeeze theorem I need its bounds which I have not been able to figure it out

OpenStudy (anonymous):

n^n = n*n*n*n*n (n times) n! = 1*2*3*...*n So n^n/n! can be written as n/1 * n/2 * n/3*...*n/n Which is n multiplied by a number of things which are all greater than or equal to 1. Since n goes to infinity, the product must go to infinity.

OpenStudy (anonymous):

Oh I know.

OpenStudy (anonymous):

Just use n. n is always less than the sequence, but approaches infinity.

OpenStudy (anonymous):

And an upper bound that works is n^n, which obviously approaches infinity.

OpenStudy (anonymous):

BAM. SANDWHICHED.

OpenStudy (anonymous):

And which is the lower bound here

OpenStudy (anonymous):

n is the lower bound. Do you see why? I rewrote the function as n/1*n/2*n/3*...*n/n which is just n multiplied by a bunch of other things which are all greater than or equal to 1. So the function is greater than n.

OpenStudy (anonymous):

n < n^n/n!

OpenStudy (anonymous):

lim n->infinity = infinity

OpenStudy (anonymous):

The upper bound is kind of extraneous after that, but n^n works well.

OpenStudy (anonymous):

Why n^n is the upper bound here?

OpenStudy (anonymous):

Because the sequence is n^n / n! n! is greater than one, so obviously if you don't divide by n!, then the sequence will be greater.

OpenStudy (anonymous):

Did I explain the lower bound part well enough? Did that make sense?

OpenStudy (anonymous):

I am still kinda confused with the lower bound here

OpenStudy (anonymous):

http://www.twiddla.com/589333

OpenStudy (anonymous):

I am multiplying n by the numbers equal or less than n so ..........

OpenStudy (anonymous):

I don't have microphone here so I will be in touch with u only through chat, will that be fine?